Consider a simple economy with two goods, shelter and food, and two consumers Brian and Davina. Brian is endowed with 200 units of shelter and 200 units of food. Davina is endowed with 100 units of shelter and 300 units of food. Suppose that Brian’s optimal demand for shelter is S = 100. Davina’s optimal demand for shelter is S = 50 + 150PF / PS where PS is the price of shelter and PF the price of food.

a. Suppose that PS = 2 and PF = 1. Draw the Edgeworth box, initial allocation and budget constraint.

Do these prices produce general equilibrium? (Yes or No)?

b. Find prices that do produce general equilibrium. (Round your answer to two decimal places (e.g., 32.16).)

PF / PC =?
Answer #1

Shelter for D 500 Equilibrium of B 300 General Equilibrium Equilibrium of D Food for B Food for D Endowment 200 -100 200 Shel

as Individual demand for sheltere: OSO + 1 CS sb100 SD - 450 + 150 lei Market demand: s= 8+ so . C: S = 150 + 150 Chapte For